Wheels Pty(Ltd) is a large private company that specialises in refurbishing old tyres. The management plans to raise equity capital. The company wants to list on the JSE and plans to raise R10 billion. The company has 25 million shares outstanding and can issue 25 million more in an offer for subscription. A bank evaluated the value of the company and estimates that a fair value for one share is R1000. If the company takes under-pricing of 20% into account, how many shares would the company have to issue to raise R10 billion through their listing?
1
Expert's answer
2020-08-17T18:17:50-0400

share price as per bank evaluation = R 1000

company takes it is underpriced by 20%

so, fair value of share actually = "\frac{1000}{100-20}"

= R 1250

no. of share to be isuued = "\frac{10 billion}{R1250}"

= "\frac{10(1000)million}{1250}"


= 8 million

so, 8 million shares have to be issued by the company.